(x^-3)^-2(x^-1y)^2/(xy^-2)^2
Answer by stanbon(48535)

(x^-3)^-2(x^-1y)^2/(xy^-2)^2
=(x^6)(x^-2y^2) / (x^2y^-4)
= (x^4y^2)/(x^2/y^4)
= x^2y^6
====================
Cheers,
Stan H.